OpenStudy (anonymous):

Solve the following using the quadratic formula. Round to the nearest tenth, if necessary. x2 + 5x - 2 = 0 {-2, 5} {-5.4, 0.4} {-1, 0.5} No real solution

OpenStudy (cwrw238):

use the quadratic formula x =[ -b +- sqrt(b^2 - 4ac)] / 2a here a = 1 , b = -5 and c = -2

OpenStudy (campbell_st):

a = 1, b = 5 and c - 2 then \[x = \frac{-5\pm \sqrt{5^2 - 4\times1\times(-2)}}{2\times1}\] just evaluate

OpenStudy (anonymous):

would it be B? o:

OpenStudy (campbell_st):

B looks like the answer

OpenStudy (anonymous):

Thanks for the help(: <3